OpenStudy (anonymous):

Find the following combinations and composites: f(x)= x=5 g(x)=[x^2] - 5x a) (f=g)(x) There's more but I just need the jist to know what to do idk how please explain ;o

OpenStudy (anonymous):

is that equal sign suppose to be a +

OpenStudy (anonymous):

whoops! yes

OpenStudy (anonymous):

here is the idea my friend: (f+g)(x)=f(x)+g(x)

OpenStudy (anonymous):

Is it really that easy?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

alright, what about this: \[\left( f | g \right)(x)\]

OpenStudy (anonymous):

is that suppose to be division sign?

OpenStudy (anonymous):

idk how to make the little circle sign, it has (f o G)(x) ??

OpenStudy (anonymous):

oh, if you wanted a circle you should have out an o there of somehting, that line is not even close to a circle

OpenStudy (anonymous):

\[f o g(x)=f(g(x))\]

OpenStudy (anonymous):

lol i know, but it is shown like that line too

OpenStudy (anonymous):

your awesome. thanks

OpenStudy (anonymous):

no problem